# NP-LoRA: Null Space Projection Unifies Subject and Style in LoRA Fusion

### v1

@@ Abstract
Low-Rank Adaptation (LoRA) fusion has emerged as a key technique for reusing and composing learned subject and style representations for controllable generation without costly retraining. However, existing methods rely on weight-based merging, where one LoRA often dominates the other, leading to interference and degraded fidelity. This interference is structural: separately trained LoRAs occupy low-rank high-dimensional subspaces, leading to non-orthogonal and overlapping representations. In this work, we analyze the internal structure of LoRAs and find their generative behavior is dominated by a few principal directions in the low-rank subspace, which should remain free from interference during fusion. To achieve this, we propose Null Space Projection LoRA (NP-LoRA), a projection-based framework for LoRA fusion that enforces subspace separation to prevent structural interference among principal directions. Specifically, we first extract principal style directions via singular value decomposition (SVD) and then project the subject LoRA into its orthogonal null space. Furthermore, we introduce a soft projection mechanism that enables smooth control over the trade-off between subject fidelity and style consistency. Experiments show NP-LoRA consistently improves fusion quality over strong baselines (e.g., DINO and CLIP-based metrics, with human and LLM preference scores), and applies broadly across backbones and LoRA pairs without retraining.
@@
Figure 1: Illustration of our motivation. (a) This task is to combine a content LoRA capturing subject identity with a style LoRA encoding artistic appearance, enabling the reuse and composition of learned generative knowledge. (b) Weighted addition of two LoRAs destroys style features in the merged results. Independently trained LoRAs, sharing the pretrained diffusion feature space, tend to occupy correlated low-rank high-dimensional subspaces (e.g., texture or color), which causes representational interference that degrades the fidelity of generated content. (c) The content LoRA is projected onto the null space of style lora, effectively avoiding overlap and preserving stylistic features.
@@
  • We provide a theoretical analysis showing that weight-based merging inherently introduces interference, and leverage SVD to identify the style subspace and construct its null space for interference-free fusion.
@@ II Related Work
Customization of diffusion models. Diffusion models [ddpm, ddim_1, ddim_2, ldm] have fueled extensive research on few-shot personalization. Approaches such as Textual Inversion [textinver_1, textinver_2, textinver_3] and DreamBooth [dreambooth] respectively learn token embeddings or fine-tune model parameters to capture subject identity. Subsequent methods, including Custom Diffusion [customdiffusion] and several training-free variants [avrahami2023break, shi2024instantbooth, xiao2025fastcomposer, xie2023smartbrush], improve efficiency by reusing or selectively updating pretrained components. Despite these advances, most approaches remain computationally demanding and struggle to generalize. Low-Rank Adaptation (LoRA) [hu2022lora, hayou2024lora+, kopiczko2023vera, ren2024melora, zhang2023lora, zhou2024lora, zi2023delta, Roy_2025_ICCV] offers an efficient alternative through rank-decomposed adapters that balance flexibility and efficiency. Among its extensions, B-LoRA [blora] learns disentangled style and content LoRAs for later fusion. In this work, we study how independently trained LoRAs can be effectively merged to achieve harmonious concept fusion.
Merging multiple LoRAs for image generation. Unlike joint training methods that learn content and style from scratch, we focus on fusing independently trained LoRAs. In image generation, LoRA merging generally falls into two categories: object composition [dong2024continually, gu2023mix, jiang2025mc, liu2023cones, yang2024lora, zhong2024multi, zhang2025rethinking] and content-style fusion [ziplora, klora, lorarar]. We target the latter, aiming to combine content and style LoRAs while preserving both subject fidelity and stylistic consistency. Among recent methods, ZipLoRA [ziplora] and K-LoRA [klora] go beyond uniform addition but still rely on weighted combinations; i.e., ZipLoRA optimizes a fusion ratio to balance content and style, while K-LoRA adaptively switches between them during diffusion. LoRA.rar [lorarar] further employs a hypernetwork trained on content-style LoRA pairs to generalize to unseen combinations. Despite these advances, existing methods fuse LoRAs through weight-based operations, implicitly assuming that the two can be merged without interference, i.e., an assumption our analysis later disproves.
@@ III Background
A fundamental requirement is that the merged LoRA should preserve the distinctive characteristics of both the content and style LoRAs without mutual interference. To satisfy this requirement, it is essential to understand how LoRA encodes content and style information internally. Therefore, we delve into the internal structure of LoRA to identify the key information that must be preserved to achieve interference-free fusion.
@@ V-A Experiment Setup
Experimental details. All experiments are conducted on the SDXL v1.0 [sdxl] and FLUX [flux] models. For our method, we set the rank to 8 and preserve all principal directions, as described in Sec. IV-A. The key hyperparameter $\mu$ in the soft projection (Eq. 21) is fixed to 0.5, with a detailed analysis provided in Sec. V-B. Implementation details are provided in the supplemental material. Experiments based on SDXL are performed on an NVIDIA RTX 3090 GPU, while those on FLUX are conducted on an NVIDIA A800 GPU.
@@ V-B Ablation on Projection Strength
Figure 8: Qualitative results on flux for generalization validation
@@
B-LoRA [blora] 0.72 0.55 0.51 0.26 0.51 0.44
ZipLoRA [ziplora] 0.75 0.56 0.55 0.24 0.52 0.44
K-LoRA [klora] 0.76 0.55 0.57 0.22 0.52 0.42
LoRA.rar [lorarar] 0.70 0.53 0.50 0.30 0.51 0.46
@@
B-LoRA [blora] 9.06% 9.38%
ZipLoRA [ziplora] 10.00% 9.38%
K-LoRA [klora] 12.19% 12.50%
LoRA.rar [lorarar] 11.25% 9.38%
@@
TABLE III: Efficiency comparison of different LoRA merging strategies. Merge (s) includes model loading time; Gen./Img. (s) is the average generation time per image.
@@
ZipLoRA [ziplora] >10>10 min 29.931
K-LoRA [klora] 12.250 60.388
@@
V-C Results
Quantitative comparisons.
In order to ensure a fair evaluation, all experiments at this stage are conducted using SDXL. Following prior works including K-LoRA [klora] (18 pairs) and ZipLoRA [ziplora] (32 pairs), we evaluate our NP-LoRA on 32 randomly selected content-style LoRA pairs, covering 5 animal subjects, 5 object-level categories, and 15 distinct artistic styles. Each pair contains 10 generated images. For the evaluation, following [ziplora, klora], we adopt CLIP [clip] and DINO [dino] similarities for jointly evaluating both subject preservation (content) and style alignment (style) [jiang2024clipdinovisualencoders, ziplora, klora]. Specifically, for CLIP we compute $S^{\text{CLIP}}_{\text{content}}$ and $S^{\text{CLIP}}_{\text{style}}$, and for DINO we compute $S^{\text{DINO}}_{\text{content}}$ and $S^{\text{DINO}}_{\text{style}}$. To obtain unified content-style similarity measures, we report both the arithmetic mean and the harmonic mean following the common F1-style aggregation scheme [van1979information, f1, ristani2016mtmc, xian2017cvpr]. Given the four similarity scores $S_{i}\in\{S^{\text{CLIP}}_{\text{content}},S^{\text{CLIP}}_{\text{style}},S^{\text{DINO}}_{\text{content}},S^{\text{DINO}}_{\text{style}}\}$, we report both the arithmetic mean $S_{\text{arith}}=\tfrac{1}{4}\sum_{i}S_{i}$ and the harmonic mean $S_{\text{harm}}=\tfrac{4}{\sum_{i}(1/S_{i})}$, where $S_{\text{arith}}$ reflects the overall trend, while the harmonic mean $S_{\text{harm}}$ emphasizes joint fidelity to both content and style. As shown in Tab. I, our method achieves the highest $S_{\text{arith}}$ and $S_{\text{overall}}$, showing its superior ability to balance subject fidelity and style consistency.
To further assess perceptual quality, we conducted a user study with 20 participants following K-LoRA [klora]. Each case included outputs from SOTA baselines and our method, along with reference subject and style images. Participants selected the result best balancing subject fidelity and style consistency. As shown in Tab. II, our method was preferred in $49.53\%$ of cases. We further employed GPT-5-based automatic evaluation, which yielded a consistent $53.13\%$ preference for our method.
@@ V-C Results
To evaluate efficiency, we compare both the merging and generation times in Tab. III. LoRA.rar [lorarar] is excluded because it depends on pretrained models for parameter fusion, and B-LoRA [blora] is also omitted since it requires training from scratch instead of merging pretrained LoRAs. ZipLoRA [ziplora] is the slowest due to additional fusion optimization, whereas K-LoRA [klora] is efficient during initialization but slows down because of dynamic fusion in the sampling stage. In contrast, NP-LoRA maintains merging and inference efficiency comparable to the Direct Merge baseline, yet consistently delivers superior results.
@@ V-D Projection Direction: Content into Style
V-E Evaluation on Flux for Generalization
We further verify the generality of our method on the FLUX [flux] backbone, as illustrated in Fig. 8. Among existing approaches, only K-LoRA provides an official implementation for FLUX, and we include it as a reference. We use publicly available LoRAs to perform content-style fusion across several pairs. While all methods produce visually plausible results on FLUX, subtle differences remain: direct merging produces entangled outputs, K-LoRA favors content, and our NP-LoRA achieves a balanced fusion of content and style. These observations demonstrate that our projection-based design generalizes well across different diffusion backbones.
@@ VI Conclusions
In this paper, we propose NP-LoRA, a merging strategy designed to harmoniously unify style and content LoRAs while avoiding destructive interference. Our method first identifies the style-critical subspace via singular value decomposition (SVD) and projects the content LoRA onto its orthogonal complement through a hard projection. We then softly attenuate the overlapping components according to the desired level of flexibility. Furthermore, we theoretically analyze why conventional weight merging fails and show why null space projection effectively resolves this issue. Comprehensive experiments, including quantitative metrics, user studies, and GPT-5 evaluations, confirm the effectiveness and preference of our method. Notably, NP-LoRA delivers significant performance gains while maintaining comparable runtime efficiency to direct merging.
Limitations. Our method builds a single-layer projection mechanism under the assumption that LoRA representations are cross-layer independent, which is shared with prior merging approaches [ziplora, klora, lorarar]. This assumption may not fully capture complex non-linear or cross-layer dependencies. Moreover, the attenuation coefficient $\mu$ in NP-LoRA is empirically chosen; future work will explore adaptive or data-driven strategies for automatic selection.
@@
This supplementary material provides additional technical details, derivations, and extended analyses to support the findings presented in the main paper. Specifically, we first derive the formulation of the proposed soft projection operator in Sec. VII and present details for efficient implementation in Sec. VIII.  Sec. IX investigates which null-space construction (U-space or V-space) better preserves stylistic fidelity.  Sec. XI provides an additional comparison with joint training to further validate the effectiveness of NP-LoRA over training-based approaches.  Sec. XII demonstrates the controllability of our method under various prompt conditions, while Sec. XIII evaluates its robustness across different random seeds. Sec. XIV extends our validation to the Flux model to assess cross-model generalization. Finally, Sec. XV details the GPT-5 evaluation protocol used for quantitative preference assessment.
@@ VII Derivation of the Soft Projection Operator
VIII Implementation Details
@@
IX Which Null-Space Construction Preserves Style: $U$-space or $V$-space?
@@
X Additional Ablation on Soft Projection
@@
XI Additional Comparison with Joint Training
@@
XII Prompt Control
We conduct experiments to assess whether our method can alter the object’s actions, the surrounding environment, or introduce new elements by adjusting the prompts. As shown in 11, after modifying the prompts, our method effectively preserves the original object’s characteristics and stylistic attributes, while seamlessly integrating new elements or scene details.
@@ XIII Robustness Test
XIV Further Generalization Validation on FLUX
To further validate the generality of NP-LoRA, we apply our method to the FLUX backbone using publicly available LoRAs covering diverse subjects and styles. As shown in Fig. 13, 14, NP-LoRA consistently preserves the balance between subject identity and stylistic fidelity across various domains. These results demonstrate that our projection-based fusion mechanism remains effective even under different diffusion architectures, highlighting the robustness and general applicability of the proposed approach.
XV Details for GPT-5 Evaluation Protocol

### v2

@@ Abstract
Low-Rank Adaptation (LoRA) fusion enables the composition of learned subject and style representations for controllable generation without retraining. However, existing methods rely on weight-based merging within a shared adaptation space, where independently trained LoRAs interfere and degrade fidelity. We show that this interference is fundamentally geometric: content and style LoRAs occupy overlapping, non-orthogonal low-rank subspaces, making weight-based fusion inherently flawed. Analyzing LoRA internal structure, we find that generative behavior is dominated by a few principal directions that must be preserved during fusion. Based on this insight, we reformulate LoRA fusion as a null-space projection problem and propose Null Space Projection LoRA (NP-LoRA), a projection-based framework that enforces subspace separation by construction. NP-LoRA extracts principal style directions via singular value decomposition (SVD) and projects the subject LoRA into the orthogonal complement of the style subspace, preventing interference. We further introduce a soft projection mechanism that provides continuous control over the trade-off between subject fidelity and style preservation. Experiments show that NP-LoRA consistently outperforms strong baselines and generalizes well across pretrained LoRA pairs without retraining.
Figure 1: Illustration of our motivation. (a) The task is to compose a content LoRA encoding subject identity with a style LoRA capturing artistic appearance for reusable and controllable generation. (b) Weighted LoRA merging degrades style fidelity due to interference: independently trained LoRAs occupy correlated, non-orthogonal low-rank subspaces in the shared diffusion feature space. (c) NP-LoRA projects the content LoRA into the null space of the style LoRA, eliminating subspace overlap and preserving stylistic features.
@@
  • Our work applies SVD to explicitly resolve subspace overlap in LoRA merging by formulating LoRA fusion as a null-space projection problem. This approach removes interfering components by projection, enabling interference-free composition of independently trained LoRAs.
@@ II Related Work
Customization and LoRA merging in diffusion models. Diffusion models [ddpm, ddim_1, ddim_2, ldm] have enabled extensive research on few-shot personalization. Early methods such as Textual Inversion [textinver_1, textinver_2, textinver_3] and DreamBooth [dreambooth] capture subject identity by learning token embeddings or fine-tuning model parameters, while later approaches, including Custom Diffusion [customdiffusion] and several training-free variants [avrahami2023break, shi2024instantbooth, xiao2025fastcomposer, xie2023smartbrush], improve efficiency by reusing or selectively updating pretrained components. Low-Rank Adaptation (LoRA) [hu2022lora, hayou2024lora+, kopiczko2023vera, ren2024melora, zhang2023lora, zhou2024lora, zi2023delta, Roy_2025_ICCV] provides a more efficient alternative via rank-decomposed adapters. Extensions such as B-LoRA [blora] disentangle style and content for later fusion, while CSGO [xing2024csgo] composes content and style in a manner similar to image style transfer. Rather than jointly learning content and style, we focus on merging independently trained LoRAs. Existing LoRA merging methods for image generation can be broadly categorized into object composition [dong2024continually, gu2023mix, jiang2025mc, liu2023cones, yang2024lora, zhong2024multi, zhang2025rethinking] and content-style fusion [ziplora, klora, lorarar, estlora]. Recent methods such as ZipLoRA [ziplora], K-LoRA [klora], and LoRA.rar [lorarar] move beyond naive addition through weighted fusion, adaptive switching, or hypernetworks. However, they all rely on weight-based merging and implicitly assume minimal interference between LoRAs, which our analysis shows to be flawed.
Orthogonality and Subspace Methods in Model Adaptation. Orthogonality is widely used in machine learning to reduce interference, often analyzed via singular value decomposition (SVD). Recent works [knots, subzero, prolora, fangalphaedit, zhang2025lion] incorporate SVD or orthogonal designs in different adaptation settings and address distinct problems. KnOTS [knots] uses SVD to extract shared bases for LoRA alignment and aggregation, but does not explicitly remove overlapping subspaces. SubZero [subzero] employs orthogonal cross-attention to mitigate content-style leakage during denoising, operating at the feature level rather than on LoRA parameters. LiON-LoRA [zhang2025lion] focuses on coordinating spatial–temporal LoRAs in video diffusion, rather than explicitly enforcing orthogonality via SVD-based projection. ProLoRA [prolora] and LoRA-Null [tang2025lora] both exploit subspace or null-space properties of LoRA, but focus on LoRA re-adaptation or initialization rather than LoRA fusion. Although these methods leverage orthogonality or SVD, they target different scenarios and objectives. In contrast, we focus on LoRA merging by formulating LoRA fusion as a null-space projection problem and applying SVD to remove overlapping components by construction, enabling interference-free composition of independently trained LoRAs.
@@ III Background
To achieve interference-free fusion, the merged LoRA must preserve the distinct characteristics of both content and style, motivating an analysis of how LoRA internally encodes and represents this information.
@@ V-A Experiment Setup
Experimental details. All experiments are conducted on the SDXL v1.0 [sdxl]. For training the LoRAs to be merged, including the training prompts and ranks, we follow the settings in K-LoRA [klora, dreambooth]. The key hyperparameter $\mu$ in the soft projection (Eq. 21) is fixed to 0.5, with a detailed analysis provided in Sec. V-B. Implementation details are provided in the supplemental material. Experiments based on SDXL are performed on an NVIDIA RTX 3090 GPU.
@@ V-B Ablation on Projection Strength
V-C Results
Quantitative comparisons.
In order to ensure a fair evaluation, all experiments at this stage are conducted using SDXL. Following prior works including K-LoRA [klora] (18 pairs) and ZipLoRA [ziplora] (32 pairs), we evaluate our NP-LoRA on 32 randomly selected content-style LoRA pairs, covering 5 animal subjects, 5 object-level categories, and 15 distinct artistic styles. Each pair contains 10 generated images, which will be clarified in Supplementary Material. For evaluation, following prior work [ziplora, klora], we adopt CLIP [clip] and DINO [dino] similarity metrics to jointly assess subject preservation (content) and style alignment (style) [jiang2024clipdinovisualencoders, ziplora, klora]. These metrics are widely used in the literature on LoRA merging.Specifically, for CLIP we compute $S^{\text{CLIP}}_{\text{content}}$ and $S^{\text{CLIP}}_{\text{style}}$, and for DINO we compute $S^{\text{DINO}}_{\text{content}}$ and $S^{\text{DINO}}_{\text{style}}$. To obtain unified content-style similarity measures, we report both the arithmetic mean and the harmonic mean following the common F1-style aggregation scheme [van1979information, f1, ristani2016mtmc, xian2017cvpr]. Given the four similarity scores $S_{i}\in\{S^{\text{CLIP}}_{\text{content}},S^{\text{CLIP}}_{\text{style}},S^{\text{DINO}}_{\text{content}},S^{\text{DINO}}_{\text{style}}\}$, we report both the arithmetic mean $S_{\text{arith}}=\tfrac{1}{4}\sum_{i}S_{i}$ and the harmonic mean $S_{\text{harm}}=\tfrac{4}{\sum_{i}(1/S_{i})}$, where $S_{\text{arith}}$ reflects the overall trend, while the harmonic mean $S_{\text{harm}}$ emphasizes joint fidelity to both content and style. As shown in Tab. I, our method achieves the highest $S_{\text{arith}}$ and $S_{\text{overall}}$, showing its superior ability to balance subject fidelity and style consistency.
@@ V-C Results
B-LoRA 0.72 0.55 0.51 0.26 0.51 0.44
ZipLoRA 0.75 0.56 0.55 0.24 0.52 0.44
K-LoRA 0.76 0.55 0.57 0.22 0.52 0.42
LoRA.rar 0.70 0.53 0.50 0.30 0.51 0.46
@@
B-LoRA 9.06% 9.38%
ZipLoRA 10.00% 9.38%
K-LoRA 12.19% 12.50%
LoRA.rar 11.25% 9.38%
@@
Following K-LoRA [klora], we further evaluated perceptual quality through a user study involving 20 participants. The study was conducted on images generated from 32 merged LoRA pairs (Tab. I), with 32 images sampled from each pair. Each case included outputs from SOTA baselines and our method, along with reference subject and style images. Participants selected the result best balancing subject fidelity and style consistency. As shown in Tab. II, our method was preferred in $49.53\%$ of cases. We further adopted a GPT-5-based automatic evaluation protocol, which is more recent than the MLLM evaluator used in [lorarar]. The results show a consistent $53.13\%$ preference in favor of our method.
TABLE III: Efficiency comparison of different LoRA merging strategies. Merge (s) includes model loading time; Gen./Img. (s) is the average generation time per image. LoRA.rar [lorarar] is excluded for its needs pretrained models, and B-LoRA [blora] is omitted for requiring scratch training.
@@
ZipLoRA >10>10 min 29.931
K-LoRA 12.250 60.388
@@
To evaluate efficiency, we compare both the merging and generation times in Tab. III. LoRA.rar [lorarar] is excluded because it depends on pretrained models for parameter fusion, and B-LoRA [blora] is also omitted since it requires training from scratch instead of merging pretrained LoRAs. ZipLoRA [ziplora] is the slowest due to additional fusion optimization, whereas K-LoRA [klora] is efficient during initialization but slows down for dynamic fusion in the sampling stage. In contrast, NP-LoRA maintains merging and inference efficiency comparable to Direct Merge baseline, yet consistently delivers superior results.
@@ VI Conclusions
In this paper, we propose NP-LoRA, a merging strategy that harmoniously unifies style and content LoRAs while avoiding destructive interference. NP-LoRA identifies a style-critical subspace via SVD and projects the content LoRA onto its orthogonal complement, followed by soft attenuation of overlapping components to control flexibility. We further provide theoretical analysis explaining the failure of conventional weight merging and the effectiveness of null-space projection. Extensive experiments, including quantitative metrics, user studies, and GPT-5 evaluations, demonstrate the effectiveness of NP-LoRA, which achieves significant performance gains with runtime efficiency comparable to direct merging.
Limitations. Our method assumes cross-layer independence of LoRA representations, as in prior merging approaches [ziplora, klora, lorarar], which may miss complex cross-layer or non-linear dependencies. In addition, the attenuation coefficient $\mu$ in NP-LoRA is empirically set, and we leave adaptive or data-driven selection for future work.
@@
This supplementary material provides additional technical details, derivations, and extended analyses to support the findings presented in the main paper. Specifically, we first derive the formulation of the proposed soft projection operator in Sec. VII and present details for efficient implementation in Sec. IX.  Sec. X investigates which null-space construction (U-space or V-space) better preserves stylistic fidelity.  Sec. XII provides an additional comparison with joint training to further validate the effectiveness of NP-LoRA over training-based approaches.  Sec. XIII demonstrates the controllability of our method under various prompt conditions, while Sec. XIV evaluates its robustness across different random seeds. Sec. XV extends our validation to the Flux model to assess cross-model generalization. Finally, Sec. XVI details the GPT-5 evaluation protocol used for quantitative preference assessment.
@@ VII Derivation of the Soft Projection Operator
VIII Randomly Sampling LoRA Pairs.
Following prior works including K-LoRA [klora] (18 pairs) and ZipLoRA [ziplora] (32 pairs), we evaluate our NP-LoRA on 32 randomly sampled content–style LoRA pairs, covering 5 animal subjects, 5 object-level categories, and 15 distinct artistic styles.
The specific content–style combinations are listed below:
dog in watercolor painting style3, in kid crayon drawing style, in oil painting style3, in watercolor painting style2
cat2 in watercolor painting style3, in oil painting style1, in oil painting style2
dog2 in sticker style, in watercolor painting style3, in oil painting style3
vase in line drawing style, in oil painting style3, in sticker style
candle in 3d rendering style2, in kid crayon drawing style, in watercolor painting style2
colorful_sneaker in watercolor painting style1, in line drawing style, in wooden sculpture
bear_plushie in oil painting style2, in wooden sculpture, in 3d rendering style
dog5 in cartoon line drawing style, in glowing 3d rendering style
wolf_plushie in flat cartoon illustration style, in abstract rainbow colored flowing smoke wave design, in 3d rendering style
dog8 in cartoon line drawing style, in flat cartoon illustration style, in glowing 3d rendering style, in sticker style
[…]
The naming of content categories and artistic styles strictly follows the conventions used in publicly released codebases, consistent with K-LoRA [dreambooth, styledrop]. Readers may refer to the corresponding open-source implementations for qualitative visualizations of each content–style pair.
IX Implementation Details
@@
X Which Null-Space Construction Preserves Style: $U$-space or $V$-space?
@@
XI Additional Ablation on Soft Projection
@@
XII Additional Comparison with Joint Training
@@
XIII Prompt Control
We conduct experiments to assess whether our method can alter the object’s actions, the surrounding environment, or introduce new elements by adjusting the prompts. As shown in 10, after modifying the prompts, our method effectively preserves the original object’s characteristics and stylistic attributes, while seamlessly integrating new elements or scene details.
@@ XIV Robustness Test
XV Further Generalization Validation on FLUX
To further validate the generality of NP-LoRA, we apply our method to the FLUX backbone using publicly available LoRAs covering diverse subjects and styles. As shown in Fig. 12, 13, NP-LoRA consistently preserves the balance between subject identity and stylistic fidelity across various domains. These results demonstrate that our projection-based fusion mechanism remains effective even under different diffusion architectures, highlighting the robustness and general applicability of the proposed approach.
Prompt used for GPT-5 evaluation.
Task: Evaluate and select the candidate image that best balances Subject Fidelity and Style Consistency.
Inputs: - First image = A.jpg (Subject Fidelity reference): core subject content to preserve. - Second image = B.jpg (Style reference): artistic style, colors, lighting to apply. - Next six images are candidates, named [Model_1]..[Model_6] in the given order.
Candidate mapping (fixed): [Model_1] = Direct [Model_2] = Klora [Model_3] = LoRArar [Model_4] = B-LoRA [Model_5] = ziplora [Model_6] = NP-LoRA
Evaluation Criteria: 1) Subject Fidelity: retain key characteristics/structure/recognizability from A.jpg. 2) Style Consistency: apply style/texture/colors/lighting present in B.jpg.
Crucial Requirement: Do not choose images that only retain the subject without the style, or only mimic the style with distorted subject. Select the BEST BALANCE.
Output Rules (IMPORTANT): You MUST respond with ONLY ONE of the following exact strings (case-sensitive): Direct Klora LoRArar B-LoRA ziplora NP-LoRA
XVI Details for GPT-5 Evaluation Protocol